Lawanda and Betty both drive the same distance to the shore. By driving 15 miles per hour faster than Betty, Lawanda can get there in 3 hours while Betty takes 4 hours. How fast does each drive?
Answer by ptaylor(1894)


Let x= Betty's speed
Then x+15=Lawanda's speed

Distance(d)=rate(r) times time(t) or d=rt
Distance Betty drives=4x
Distance Lawanda drives=3(x+15)
Now we are told that they both drive the same distance. So our equation to solve is:
4x=3(x+15) get rid of parens
4x=3x+45 subtract 3x from both sides
4x-3x=3x-3x+45 collect like terms
x=45 mph--------------------------------------Betty's speed
x+15=45+15=60 mph ------------------------Lawanda's speed
CK
3*60=4*45
180=180